A 53.7kg spacewalking astronaut pushes off a 637.0kg satellite, exerting a 120.0N force for the 0.540s it takes him to straighten his arms. How far apart are the astronaut and the satellite after 6.72min?
Answer #1

a ast = F/m = 120N/ 53.7 kg = 2.23 m/s^2

from the kinamtic equaiton

vf = u+ at = 0 + 2.23( 0.540) = 1.2 m/s

x = vf t = 1.2( 6.72 ( 60) = 486.54 m

a sat = -120/637 = -0.188 m/s^2

v = u+ at = 0 + -0.188 ( 0.540) =-0.1017 m/s

x = vf t = -0.1017 m/s( 6.72 ( 60) = -41 m

separation x = xa- xs = 486.54- ( -41) = 527.55 m